OpenStudy (cutestgirlaround):

−1/8(2h+4)+1/2=−7/4 H= ??? A. 7 B. 25/16 C. 7/16 D. 25

OpenStudy (cutestgirlaround):

@Awolflover1

Awolflover1 (awolflover1):

\[\frac{ -1 }{ 4 } h=\frac{ -7 }{ 4 }\] remember Combine Like Terms? and Distribute?

OpenStudy (cutestgirlaround):

yeah I know but for some reason I cant get it I think 7/16 but I figured I would ask

Awolflover1 (awolflover1):

hmm did you Multiply both sides by -4/1 ?

OpenStudy (cutestgirlaround):

yes

OpenStudy (cutestgirlaround):

but maybe I did it wrong

Awolflover1 (awolflover1):

\[( \frac{ -4 }{ 1 } )*( \frac{ -1 }{ 4 } h)=( \frac{- 4 }{ 1 } )*( \frac{ -7 }{ 4 } )\]

OpenStudy (cutestgirlaround):

so its 4h=7?

Awolflover1 (awolflover1):

there you go

OpenStudy (cutestgirlaround):

thanks again I think i'm getting a hang of this